2x+y-z=8


-3x-y+2z=-11


-2x+y+2z=-3


1
Expert's answer
2022-03-15T12:06:47-0400

"2x+y-z=8\\
-3x-y+2z=-11\\
-2x+y+2z=-3\\
\begin{pmatrix}
 2 & 1&-1&|&8 \\
 -3 & -1&2&|&-11\\
-2&1&2&|&-3
\end{pmatrix}\to"

IIr+Ir("\frac{3}{2}" )

IIIr+Ir

"\to\begin{pmatrix}
 2 & 1&-1&|&8 \\
 0& \frac{1}{2}&\frac{1}{2}&|&1\\
0&2&1&|&5
\end{pmatrix}\to"

IIr(2)

IIIr+IIr(-2)

"\to\begin{pmatrix}
 2 & 1&-1&|&8 \\
 0 & 1&1&|&2\\
0&0&-1&|&1
\end{pmatrix}\to"

IIIr(-1)

IIr+IIIr(-1)

Ir+IIIr(1)

"\to\begin{pmatrix}
 2 & 1&0&|&7 \\
 0 & 1&0&|&3\\
0&0&1&|&-1
\end{pmatrix}\to"

Ir+IIr(-1)

"\to\begin{pmatrix}
 2 & 0&0&|&4 \\
 0 & 1&0&|&3\\
0&0&1&|&-1
\end{pmatrix}\to"

Ir("\frac{1}{2}" )

"\to\begin{pmatrix}
 1& 0&0&|&2 \\
 0 & 1&0&|&3\\
0&0&1&|&-1
\end{pmatrix}"

"x=2,y=3,z=-1\\(2,3,-1)"